OpenStudy (anonymous):

Maria and Ming start at the same point and drive in opposite directions. Maria drives 38 mph and Ming drives 38 mph. How far apart will they be in 2 hours? THIS IS AN EASY SCIENCE QUESTION, BUT I NEED HELP WILL GIVE MEDAL!!!

OpenStudy (shadeyghost):

it's more of a math question, but i can help

OpenStudy (shadeyghost):

ok, so they are both going 38 mph, mph being "Miles Per Hour"

OpenStudy (shadeyghost):

so since they both drive for 2 hours, all you have to do is multiply 38 by 2

OpenStudy (shadeyghost):

BUT, there is more now since they are driving in opposite directions, you multiply the answer by 2

OpenStudy (anonymous):

@oneofakind123 It is a math question really, not biology. Listen to shadeyghost

OpenStudy (shadeyghost):

so, the answer of (38 * 2) * 2

OpenStudy (anonymous):

or \[38*2^{2}\]or\[2^{3}*19\]

OpenStudy (shadeyghost):

uh, no, not really

OpenStudy (anonymous):

Uh, yes really. 2 times 2 is the same as 2 squared. 38 is =19 times 2, thus 2 times 2 times 2 times 19 or 2 cubed tiems 19.

OpenStudy (shadeyghost):

well, your 38 * 2^2 is right, but we have no way of getting that 2^3

OpenStudy (anonymous):

\[38/2=19\] Thus \[2*19=38; 2*2*(2*19)=2^{3}*19\]